OpenStudy (anonymous):

Will fan and medal! 2. (10)9-(10)11 a. -2 b. -20 c. 200 d. 8 4. 6(6)+6(4) a. 60 b. 10 c. 16 d. 12 6. Use the distributive property to multiply: 5(2t-5) a.10t-5 b. 7t-25 c. 10t-25 d. -15t

OpenStudy (anonymous):

yes don't make a big deal

OpenStudy (just_one_last_goodbye):

@Anestasia want me to help you? ^_^

OpenStudy (anonymous):

yes please

OpenStudy (camerondoherty):

For the first one \[9 \times 10=90\]\[10\times 11= 110\]Then you subract \[90-110=~?\]

OpenStudy (just_one_last_goodbye):

ok lets see for the first question (10)9-(10)11 just multiply the number out of the "( )" and with the number inside them

OpenStudy (just_one_last_goodbye):

cammi honey ur confusing the child she still hasn't replied

OpenStudy (anonymous):

Holy poop yall are really helpful!

OpenStudy (anonymous):

And its not that I'm confused, Im just waiting for her to finish so I don't freak out!

OpenStudy (just_one_last_goodbye):

lol @Anestasia shes asking what 90 - 110= ?

OpenStudy (anonymous):

You cant minus 110 from 90 tho..?

OpenStudy (camerondoherty):

For the second one \[6 \times 6 = 36\]\[6 \times 4 = 24\]\[36+24 = ?\]

OpenStudy (camerondoherty):

You can, but it becomes a negative

OpenStudy (anonymous):

2. 60!

OpenStudy (camerondoherty):

Yep!

OpenStudy (just_one_last_goodbye):

good ^_^

OpenStudy (anonymous):

So 1 is -20?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about the last one?

OpenStudy (camerondoherty):

And for the last one:  So for here you multiply as the arrows indicate \[5 \times 2t = 10t \] \[5 \times -5 = -25 \] So its \[10t - 25 \]